What is the result of mixing 25.0 mL of 0.0010 M NaCl solution with 75.0 mL of 0.0050 M AgNO3 solution? Assume volumes are additive.
Answer #1

Molarity of new solution is 0:36M__ Ag Noslat nachlag) Agilis) + ali + NaNO3- moles of AgNO3= Molarity & volume (4) 0.0050 x^ Moles of Moles of All formed= 2.5x10-s moles. AgNO3 left = 3.75*10 - 2.5x105 - 3.5810-4 moles. Moles of Nano3 formed = 2.5Please give a thumbs-up.

Do comment if you have any doubts

Thank you

Best of luck